If you were going to graph the equation -2x+y=3, construct an appropriate table of values for the equation (Do not actually graph the equation) (5 marks)

Answer by Boreal(15235)

-2x+y=3
set each one separately equal to 0 and solve for the other
x=0, y=3
x=-3/2, y=0
x=1, y=5
x=2, y=7
x=-1, y=1
x=-2, y=-1